Shift Summary for day shift, 08/11/99

Sandrine Laplace (laplace@hep.uchicago.edu)
Wed, 11 Aug 1999 15:42:18 -0500

Shift Summary for day shift, 08/11/99

On Shift: Sandrine, Ashkan

Hours of beam delivered at all intensities: 0
Hours of beam delivered at nominal intensity: 0
Hours of E832 data (special runs not included): 0
Hours of special runs (laser scans, muon runs...): 0

Accounting of lost time: beam recovery in progress.

========================================================================
Runs ended this shift:

14189 to 14197 of cosmics, random pulser runs.

Tapes brought to FCC this shift:

tapes from yesterday that have not been vaulted:

QKC246-248
QKS246-248
QKK246-248
KQB086

(including tapes for online split)

========================================================================
Accesses during shift: none

========================================================================
Overview of shift:

No beam, recovery in progress. Just kept running with cosmics.

========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------

14189 7:00 New crew on the boat. Cosmic run in progress.
Sidney started copying some tapes.

8:15 TeV recovering from shutdown. No estimate.

8:30 Gas Check List completed. Still no estimate
for beam.

14190 10:30 Stop cosmic run 14189. Start random pulser
run number 14190 at high rate (100 kHz) under
the request of RickK.
Kfrend at the beginning of the run.
6 spills then hangs up...
Kfrend fixes it, then runs fine.

JasonH and RickK do some tests to find out
why we need to do a kfrend at the beginning
of the runs (for 'missing FFFF' error). Discover
that if crate 53 is reinitialized AFTER the run
is initialized, then no errors at start. Rick
concludes that if initialization is moved from
stage 3 to stage 4 the problem would therefore
be fixed. Taku has been asked to look into the
possibility of doing this. For now, we'll just
have to live with kfrending after runs start.

11:00 Sydney finishes copying tapes.

12:00 Estimate for beam : one hour.

13:20 TeV injection quench. No estimate for beam.

14:55 Laser scan from last night shows that channel
1691 has very low HCC threshold (0.6 GeV).
JasonH raises HV down HV for this channel and
loads new gain constants valid from run 14197.

15:10 SachaL changed delay jumpers of DYC to 1microsec.
in ADC FERA crate 8 (it was 4 microsec. before
after it was swapped with VV' DYC from cr.51).

15:24 MC calls to ask for the desired beam intensity...
We're getting closer ! Asked for 6.8e12,
plan to start with a muons run.


========================================================================
Follow up Notes on Previous Shift

========================================================================
Items to Follow up on Next Shift

Please inform Sasha Glazov after taking the muon run.